# Convergence Analysis of Natural Gradient Descent for Over-parameterized Physics-Informed Neural Networks

### v1

@@ 7.1. Proof of Lemma 4
Therefore, applying Lemma 8 yields that with probability at least $1-\delta$,
$|\sum\limits_{r=1}^{m}\frac{1}{m}(X_{r}(ij)-\mathbb{E}[X_{r}(ij)])|\lesssim\frac{d^{2}}{\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{m}(\log(\frac{1}{\delta}))^{2},$
which means
$|\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}w_{r}}\rangle|\lesssim\frac{d^{2}}{n_{1}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{n_{1}m}(\log(\frac{1}{\delta}))^{2}.$ (58)
@@
$\left\|\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}$
@@
$\left\|\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{n_{2}m}.$
Thus applying Lemma 8 yields that with probability at least $1-\delta$,
$|\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle|\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}\log(\frac{1}{\delta})$ (59)
@@
$|\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle|\lesssim\frac{d^{2}}{n_{2}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{n_{2}m}\log(\frac{1}{\delta}).$ (60)
@@
$\lesssim\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta})+\frac{d^{4}}{m^{2}}(\log(\frac{n_{1}+n_{2}}{\delta}))^{4}$
$\lesssim\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta}).$
Thus when $\sqrt{\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta})}\lesssim\frac{\lambda_{0}}{4}$, i.e.,
$m=\Omega(\frac{d^{4}}{\lambda_{0}^{2}}\log(\frac{n_{1}+n_{2}}{\delta})),$
@@ 7.2. Proof of Lemma 5
$\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}w_{r0}(k)x_{p0}\\ w_{r0}(k)\bm{x}_{p1}-2\bm{w}_{r1}(k)\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\bm{x}_{p}.$
Similar to Lemma 4, it suffices to bound $\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}$, which can in turn allows us to bound each entry of $\bm{H}(\bm{w})-\bm{H}(0)$.
For $i\in[n_{1}]$ and $j\in[n_{1}]$, we have
$H_{ij}(\bm{w})$ $=\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}}{\partial\bm{w}_{r}},\frac{\partial s_{j}}{\partial\bm{w}_{r}}\rangle$
@@
In fact, when bounding the corresponding terms for $H_{ij}(\bm{w})-H_{ij}(0)$, the first four classes can be grouped into one category. They are of the form $f_{1}(\bm{w})f_{2}(\bm{w})f_{3}(\bm{w})f_{4}(\bm{w})$, where for each $1\leq i\leq 4$, $f_{i}(\bm{w})$ is Lipschitz continuous with respect to $\|\cdot\|_{2}$ and $|f_{i}(\bm{w})|\lesssim\|\bm{w}\|_{2}$ (Note that $\sigma^{{}^{\prime}}(\cdot)=(\sigma^{{}^{\prime\prime}}(\cdot))^{2}$). On the other hand, when $\|\bm{w}_{1}-\bm{w}_{2}\|_{2}\leq R\leq 1$, we can deduce that
@@
Thus, for the terms in $H_{ij}(w)-H_{ij}(0)$ that belong to the first four classes, we can deduce that they are less than $CR(\|\bm{w}_{r}(0)\|_{2}^{3}+1)$, where $C$ is a universal constant.
@@
where the definition of $A_{jr}$ is same as (44) for the regression problems.
Thus, we can deduce that for the terms in $H_{ij}(w)-H_{ij}(0)$ that belong to the classes $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}}$, they are less than
$C\left[(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{3}+R(\|\bm{w}_{r}(0)\|_{2}^{3}+1)\right],$
@@
$\lesssim I\{A_{ir}\lor A_{jr}\}\|\bm{w}_{r}(0)\|_{2}^{4}+R\|\bm{w}_{r}(0)\|_{2}^{3}+R.$
Combining the upper bounds for the six classes, we have that
$|H_{ij}(\bm{w})-H_{ij}(0)|$ $\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{3})+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{4}+R\right]$ (61)
$\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4})+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{4}+R\right],$
@@
$P(|w_{ri}(0)|^{2}\geq 2\log(\frac{2}{\delta}))\leq\delta$
@@
$P(\|w_{r}(0)\|_{2}^{2}\geq 2(d+2)\log(\frac{2(d+2)}{\delta}))\leq\delta,$
we obtain that
$P(\exists r\in[m],\|w_{r}(0)\|_{2}^{2}\geq 2(d+2)\log\frac{2m(d+2)}{\delta})\leq\delta.$ (62)
Let $M=2(d+2)\log(\frac{2m(d+2)}{\delta})$, then
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}$
$=\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}\leq M\}+\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}>M\}$
$\leq\frac{M^{2}}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}>M\}.$
@@
Moreover, with probability at least $1-\delta$, $I\{\|w_{r}(0)\|_{2}^{2}>M\}=0$ holds for all $r\in[m]$.
@@
For $i\in[n_{1}],j\in[n_{1}+2,n_{2}]$ and $i\in[n_{1}+1,n_{2}],j\in[n_{2}]$, from the form of $\frac{\partial h_{j}}{\partial\bm{w}_{r}}$, i.e.,
$\frac{\partial h_{j}}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{n_{2}m}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j},$
we can obtain similar results for the terms $\langle\frac{\partial s_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\rangle$ and $\langle\frac{\partial h_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\rangle$.
With all results above, we have that with probability at least $1-\delta-(n_{1}+n_{2})e^{-mR}$,
@@
7.3. Proof of Lemma 6
@@
Similar to the proof of Lemma 3, we have
$s_{p}(k+1)-s_{p}(k)$ $=\left[s_{p}(k+1)-s_{p}(k)-\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle\right]+\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (63)
@@
For the second term $I_{2}^{p}(k)$, we have that
$I_{2}^{p}(k)$ $=\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (64)
$=\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},-\eta\frac{\partial L(k)}{\partial\bm{w}}\rangle$
$=-\sum\limits_{r=1}^{m}\eta\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\frac{\partial L(k)}{\partial\bm{w}_{r}}\rangle$
$=-\sum\limits_{r=1}^{m}\eta\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\sum\limits_{t=1}^{n_{1}}s_{t}(k)\frac{\partial s_{t}(k)}{\partial\bm{w}_{r}}+\lambda\sum\limits_{j=1}^{n_{2}}h_{j}(k)\frac{\partial h_{j}(k)}{\partial\bm{w}_{r}}\rangle$
$=-\eta[\bm{H}(k)]_{p}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},$
@@
$h_{j}(k+1)-h_{j}(k)$ $=\left[h_{j}(k+1)-h_{j}(k)-\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle\right]+\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (65)
@@
$I_{2}^{n_{1}+j}(k)=-\eta[\bm{H}(k)]_{n_{1}+j}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}.$ (66)
@@
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}-\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}$ $=\bm{I}_{1}(k)+\bm{I}_{2}(k)$
$=\bm{I}_{1}(k)-\eta\bm{H}(k)\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}.$
@@
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k).$
7.4. Proof of Theorem 2
Proof.
The proof strategy is similar to that for Theorem 1. Our induction hypothesis is the following boundedness of the hidden weights and convergence rate of the empirical loss.
@@ 7.4. Proof of Theorem 2
where $B=\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}+1$.
From (62), we know that with probability at least $1-\delta$, $\|\bm{w}_{r}(0)\|_{2}\leq\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}$ holds for any $r\in[m]$. Thus, if we can prove that $\bm{w}_{r}(t)$ is closed enough to $\bm{w}_{r}(0)$, then $\|\bm{w}_{r}(t)\|_{2}\leq B$.
Corollary 2 (Lemma 4.1 in [19]).

### v3

@@ 9. Proof of Section 3
Therefore, applying Lemma D.1 yields that with probability at least $1-\delta$,
$\left|\sum\limits_{r=1}^{m}\frac{1}{m}\left[X_{r}(ij)-\mathbb{E}X_{r}(ij)\right]\right|\lesssim\frac{d^{2}}{\sqrt{m}}\sqrt{\log\left(\frac{1}{\delta}\right)}+\frac{d^{2}}{m}\left(\log\left(\frac{1}{\delta}\right)\right)^{2},$
which directly yields that
$\left|\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{p}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{p}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]\right|\lesssim\frac{d^{2}}{n_{1}\sqrt{m}}\sqrt{\log\left(\frac{1}{\delta}\right)}+\frac{d^{2}}{n_{1}m}\left(\log\left(\frac{1}{\delta}\right)\right)^{2}.$ (34)
@@
$\left\|\left\langle\frac{\partial s_{i}(\bm{w}(0)}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\left\langle\frac{\partial s_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}$
@@
$\left\|\left\langle\frac{\partial h_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\left\langle\frac{\partial h_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{n_{2}m}.$
Thus applying Lemma D.1 yields that with probability at least $1-\delta$,
$\left|\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]\right|\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}\sqrt{m}}\sqrt{\log\left(\frac{1}{\delta}\right)}+\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}\log\left(\frac{1}{\delta}\right)$ (35)
@@
$\left|\sum\limits_{r=1}^{m}\left\langle\frac{\partial h_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle-\mathbb{E}_{\bm{w}(0)}\left[\sum\limits_{r=1}^{m}\left\langle\frac{\partial h_{i}(\bm{w}(0))}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w}(0))}{\partial\bm{w}_{r}}\right\rangle\right]\right|\lesssim\frac{d^{2}}{n_{2}\sqrt{m}}\sqrt{\log\left(\frac{1}{\delta}\right)}+\frac{d^{2}}{n_{2}m}\log\left(\frac{1}{\delta}\right).$ (36)
@@
$\lesssim\frac{d^{4}}{m}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)+\frac{d^{4}}{m^{2}}\left(\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)^{4}$
$\lesssim\frac{d^{4}}{m}\log\left(\frac{n_{1}+n_{2}}{\delta}\right).$
Thus when $\sqrt{\frac{d^{4}}{m}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)}\lesssim\frac{\lambda_{0}}{4}$, i.e.,
$m=\Omega\left(\frac{d^{4}}{\lambda_{0}^{2}}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right),$
@@
$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{mn_{1}}}\left[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}w_{r0}x_{p0}\\ w_{r0}\bm{x}_{p1}-2\bm{w}_{r1}\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}\right].$
It suffices to bound $\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}$, which can in turn allows us to bound each entry of $\bm{H}(\bm{w})-\bm{H}(0)$.
For $i\in[n_{1}]$ and $j\in[n_{1}]$, we have that
$H_{ij}(\bm{w})$ $=\sum\limits_{r=1}^{m}\left\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\right\rangle$
@@
In fact, when bounding the corresponding terms for $H_{ij}(\bm{w})-H_{ij}(0)$, the first four classes can be grouped into one category. They are of the form $f_{1}(\bm{w})f_{2}(\bm{w})f_{3}(\bm{w})f_{4}(\bm{w})$, where for each $i$ ($1\leq i\leq 4$), $f_{i}(\bm{w})$ is Lipschitz continuous with respect to $\|\cdot\|_{2}$ and $|f_{i}(\bm{w})|\lesssim\|\bm{w}\|_{2}$ (Note that $\sigma^{{}^{\prime}}(\cdot)=(\sigma^{{}^{\prime\prime}}(\cdot))^{2}$). On the other hand, when $\|\bm{w}_{1}-\bm{w}_{2}\|_{2}\leq R\leq 1$, we can deduce that
@@
Thus, for the terms in $H_{ij}(\bm{w})-H_{ij}(0)$ that belong to the first four classes, we can deduce that they are less than $CR(\|\bm{w}_{r}(0)\|_{2}^{3}+1)$, where $C$ is a universal constant.
@@
where the event $A_{jr}$ has been defined in (36).
Thus, we can deduce that for the terms in $H_{ij}(\bm{w})-H_{ij}(0)$ that belong to the classes $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}}$, they are less than
$C\left[(I\{A_{ir}\}+I\{A_{jr}\})(\|\bm{w}_{r}(0)\|_{2}^{3}+1)+R(\|\bm{w}_{r}(0)\|_{2}^{3}+1)\right],$
@@
$\lesssim I\{A_{ir}\lor A_{jr}\}\|\bm{w}_{r}(0)\|_{2}^{4}+R(\|\bm{w}_{r}(0)\|_{2}^{3}+1).$
Combining the upper bounds for the terms in the six classes, we have that
$|H_{ij}(\bm{w})-H_{ij}(0)|$ $\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}\left(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{3}\right)+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})(\|\bm{w}_{r}(0)\|_{2}^{4}+\|\bm{w}_{r}(0)\|_{2}^{3}+1)+R\right]$ (37)
$\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}\left(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4}\right)+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})(\|\bm{w}_{r}(0)\|_{2}^{4}+1)+R\right],$
@@
$P\left(|w_{ri}(0)|^{2}\geq 2\log\left(\frac{2}{\delta}\right)\right)\leq\delta$
@@
$P\left(\|\bm{w}_{r}(0)\|_{2}^{2}\geq 2(d+2)\log\left(\frac{2(d+2)}{\delta}\right)\right)\leq\delta.$
This implies that
$P\left(\exists r\in[m],\|\bm{w}_{r}(0)\|_{2}^{2}\geq 2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)\right)\leq\delta.$ (38)
Let $M=2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)$, then
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|\bm{w}_{r}(0)\|_{2}^{4}$
$=\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}\leq M\}+\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}$
$\leq\frac{M^{2}}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}.$
@@
Moreover, from (37), we have that with probability at least $1-\delta$, $I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}=0$ holds for all $r\in[m]$.
@@
For $i\in[n_{1}],j\in[n_{1}+2,n_{2}]$ and $i\in[n_{1}+1,n_{2}],j\in[n_{2}]$, from the form of $\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}$, i.e.,
$\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{n_{2}m}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j},$
we can obtain similar results for the terms $\left\langle\frac{\partial s_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\right\rangle$ and $\left\langle\frac{\partial h_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\right\rangle$.
With all results above, we have that with probability at least $1-\delta-n_{1}e^{-mR}$,
@@
9.3. Proof of Lemma B.1
Indeed, the stringent requirement of the learning rate in [10] stems from an inadequate decomposition method for the residual. Specifically, in [21], the decomposition for the residual in the $(k+1)$-th iteration is same as the one in [10], i.e.,
$\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}=\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}+\left[\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}-\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right],$ (39)
which leads to the requirements that $\eta=\mathcal{O}(\lambda_{0})$ and $m=Poly(n_{1},n_{2},1/\delta)$. Thus, it requires a new approach to achieve the improvements for $\eta$ and $m$. In fact, we can derive the following recursion formula.
Lemma 9.1.
For all $k\in\mathbb{N}$, we have
$\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}+\bm{I}_{1}(k),$ (40)
where
$\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n_{1}+n_{2}}(k))^{T}\in\mathbb{R}^{n_{1}+n_{2}}$
and for $p\in[n_{1}]$,
for $j\in[n_{2}]$,
In the recursion formula (39), $\bm{I}_{1}(k)$ serves as a residual term. From the proof, we can see that $\|\bm{I}_{1}(k)\|_{2}=\mathcal{O}(1/\sqrt{m})$ and thus, as $m$ becomes large enough, only the term $\bm{I}-\eta\bm{H}(k)$ is significant. This observation is the reason for the requirement of $\eta$.
@@
First, we have
$s_{p}(k+1)-s_{p}(k)$ $=\left[s_{p}(k+1)-s_{p}(k)-\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle\right]+\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle$ (43)
@@
For the second term $I_{2}^{p}(k)$, from the updating rule of gradient descent, we have that
$I_{2}^{p}(k)$ $=\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle$ (44)
$=\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},-\eta\frac{\partial L(k)}{\partial\bm{w}}\right\rangle$
$=-\sum\limits_{r=1}^{m}\eta\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\frac{\partial L(k)}{\partial\bm{w}_{r}}\right\rangle$
$=-\sum\limits_{r=1}^{m}\eta\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\sum\limits_{t=1}^{n_{1}}s_{t}(k)\frac{\partial s_{t}(k)}{\partial\bm{w}_{r}}+\sum\limits_{j=1}^{n_{2}}h_{j}(k)\frac{\partial h_{j}(k)}{\partial\bm{w}_{r}}\right\rangle$
$=-\eta\left[\sum\limits_{t=1}^{n_{1}}\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\frac{\partial s_{t}(k)}{\partial\bm{w}_{r}}\right\rangle s_{t}(k)+\sum\limits_{j=1}^{n_{2}}\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\frac{\partial h_{j}(k)}{\partial\bm{w}_{r}}\right\rangle h_{j}(k)\right]$
$=-\eta[\bm{H}(k)]_{p}\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix},$
@@
$h_{j}(k+1)-h_{j}(k)$ $=\left[h_{j}(k+1)-h_{j}(k)-\left\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle\right]+\left\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle$ (45)
@@
$I_{2}^{n_{1}+j}(k)=-\eta[\bm{H}(k)]_{n_{1}+j}\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}.$ (46)
@@
$\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}-\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}$ $=\bm{I}_{1}(k)+\bm{I}_{2}(k)$
$=\bm{I}_{1}(k)-\eta\bm{H}(k)\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}.$
@@
$\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}+\bm{I}_{1}(k),$
which is exactly (39), the new recursion formula we need to prove. ∎
9.4. Proof of Theorem 3.7
Similar to [10] and [21], we prove Theorem 3.7 by induction. Our induction hypothesis is the following convergence rate of the empirical loss and upper bounds for the weights.
@@
where $B=\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}+1$ and $L(k)$ is an abbreviation of $L(\bm{w}(k))$.
From the update formula of gradient descent, we can directly derive the following corollary, which indicates that under over-parameterization, the weights are closed to their initializations.
Corollary 9.2.